Tohban Report 2017-04-05

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 29 March 2017
End Date: 5 April 2017
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Juan Carlos Martinez
List all reports


Solar Activity

The Sun was active this week, producing seven M-class flares. Currently there are five AR on disk, but four of them are expected to rotate off the disk very soon.

How many GOES flares occurred?

 Flares above B, C, M, X class were     39    30     7     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     21    17     5     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were     12    11     2     0

There were RHESSI flares/GOES flares 188 / 76 over the time range 28-Mar-17 04-Apr-17

Memory Management

Early the week, the SSR was at 0% at the end of each day's contacts. In response to the high solar activities and the turn-on of some detectors, the SSR reached 80% and the decimation was switched to active/vigorous from normal/vigorous.

Spacecraft Status

Detectors 3 and 6 are recording events. Due to the increased solar activities, D1, 5 and 8 were turned on. D5 was turned off after ~24 hours due to its unexpectedly high event rate.

The cold tips are 129.5 and 128.0 K as of 5-Apr-2017. The cold plates, 1 and 2 are 153.3 and 151.8 K. Cryocooler power is steady at 76W. See:

for a summary plot of recent temperatures, efficiencies and accelerometer values.

Data Gaps

Relatively large (>300 sec) data gaps as follows:

GAP START TIME              GAP END TIME                   GAP (SEC)
2017-03-30T19:40:00.000 -- 2017-03-30T20:05:00.000       1500.0000
2017-04-02T06:00:00.000 -- 2017-04-02T06:25:00.000       1500.0000
2017-04-02T11:55:00.000 -- 2017-04-02T12:10:00.000       900.00000
2017-04-03T21:05:00.000 -- 2017-04-03T21:20:00.000       900.00000
2017-04-03T21:25:00.000 -- 2017-04-04T00:00:00.000       9300.0000

Detector issues

Detectors 1, 3, 6 and 8 are powered on, and are recording events. While the AR2644 that produced M-class flares is now at the limb, there is still another big AR on disk. So it was determined to keep the detectors 1 and 8 on unless there is an issue with the temperatures.


For the flare of April 3 14UT, pileup was relatively high (40% at 30 keV) with the attenuator state A1. It was decided to revisit the current thresholds of the thin/thick attenuator logic.

Spacecraft Management

Decimation Active/Vigorous
HLAT Decimation Rear decimation
Night time data (fronts) plus/minus 4 minutes from daylight
Night time data (rears) plus/minus 4 minutes from daylight
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation None
Detector problems? None
Personal tools